OpenStudy (rock_mit182):

find: h(x) ^ -1 of h(x) = 9x +13

OpenStudy (rock_mit182):

x = 9y +13 x-13 = 9y \[y = \-frac{ x }{ 9 } - \frac{ 13 }{ 9 }\]

OpenStudy (rock_mit182):

@iPwnBunnies @SolomonZelman @ganeshie8

OpenStudy (rock_mit182):

find:\[(h ° h ^{-1} )\]

OpenStudy (rock_mit182):

h ( x/9 - 13/ 9)

OpenStudy (rock_mit182):

(h ° h ^-1 ) = (9x +13)/(9) - (13)/(9)

OpenStudy (rock_mit182):

(h ° h ^-1)(5) = (9(5)+13)/(9) - 13/9

OpenStudy (rock_mit182):

anyone could check my work out ?

OpenStudy (anonymous):

Where'd the x go?

OpenStudy (anonymous):

Oh... I see... carry on :)

OpenStudy (anonymous):

I'm watching you... anxiously waiting for you to slip up... and you haven't... yet ;)

OpenStudy (rock_mit182):

(h ° h ^-1)(5 ) = 5

OpenStudy (rock_mit182):

h(x )^ -1 = x/9 - 13/9

OpenStudy (rock_mit182):

there you have, my two answers

OpenStudy (anonymous):

Very good. Now, pro-tip, if you're like me and don't like wasting precious time on things that are actually so simple, just remember the very essence of inverses... \[\Large\left( f\circ f^{-1}\right)(x) = x\] It won't even matter how complicated or hard the function is... if you compose a function with its inverse, it does NOTHING :D So, next time you see anything like it, say \[\Large \left( h\circ h^{-1}\right)(9)\] don't hesitate, the answer is 9!!!!!! ^_^

OpenStudy (rock_mit182):

lol thanks a lot dude, nice tip